# ArticuBot: Learning Universal Articulated Object Manipulation Policy via Large Scale Simulation

### v1

@@ IV-B Policy Learning with a Hierarchical Policy Representation
We then perform cross attention between the scene point cloud and the current end-effector points with Rotary Position Embedding (RoPE) [45], which generates the updated features for current end-effector points as $\{f^{\text{obs-scene}}_{i}\}_{i=1}^{4}$. We generate the features for the goal end-effector points in the same way as for the current end-effector points: $f^{goal}_{i}=[v^{goal}_{i},\text{MLP}^{goal}(ee^{goal}_{i},\delta_{i},\delta^{\prime}_{i})]$. We perform cross attention between the current end-effector points and the goal end-effector points, also with Rotary Position Embedding (RoPE) [45], which produces another set of updated features for the current end-effector points $\{f^{\text{obs-goal}}_{i}\}_{i=1}^{4}$. The final latent embedding used for diffusion is the concatenation of the above two features: $[f^{\text{obs-scene}}_{1},f^{\text{obs-goal}}_{1},...,f^{\text{obs-scene}}_{4},f^{\text{obs-goal}}_{4}]$. This latent embedding is used as the conditioning for an action generation UNet diffusion head, which takes as input this latent conditioning, a noisy version of the action, a denoising time step, and predicts the noise. At test time, we use DDIM [43] as the denoising scheduler to generate the actions.
@@ V-C Comparison of Different High-level policies
Please see Appendix -E for more details about these baselines.
@@
Replacing low-level policy with a motion planner $0.01\pm 0.0$
@@ V-D Ablation Studies
  • Replacing low-level policy with a motion planner: This ablation does not use a low-level policy for moving the robot end-effector. Instead, for the first stage of grasping the handle, we use a motion planner to reach the predicted end-effector goal pose; after grasping, for the stage of the opening the door, we use an IK controller to reach the predicted end-effector pose.
More details of these ablations can be found in Appendix -G. We compare to these ablations when training with 200 objects without camera randomizations. The results are shown in Table II. We find that using a Point Transformer, not predicting the per-point weights, or predicting a per-point goal end-effector 6D orientation instead of per-point displacements to the goal end-effector points, all lead to worse performance, supporting the effectiveness of our design choices in ArticuBot. Predicting a per-point 6D orientation and averaging them leads to a large drop in performance because it is difficult to correctly compute the average of multiple 6D orientations; in ArticuBot, we avoid this by representing the goal end-effector pose as a collection of points and averaging the displacement to these points. Replacing the low-level policy with motion planning and an IK controller results in very poor performance. We hypothesis it could be due to two reasons: First, sometimes the predicted end-effector pose for grasping the handle have small collisions with the handle, which works fine for a learned low-level policy, but motion planning often fails due to such collisions. Second, when opening the door, the solution returned by the IK controller usually tries to reach the opening end-effector pose via a trajectory that minimizes the joint angle distance, instead of following a curved trajectory that would actually open the door, which often causes the gripper to slip off the handle. This shows the importance of using a learned low-level policy.
@@ V-E Comparison with Prior Articulated Object Manipulation Methods
The results are shown in Table III, tested without camera randomizations. The grasping success rate of AO-Grasp is much lower than ArticuBot. We find that AO-Grasp often proposes grasps at the edges of the point cloud (e.g., the side wall of a drawer; see Appendix -F for visuals). This likely stems from its training data, which includes many objects in a partially opened state where edge grasps are valid. However, in our test cases, objects are usually closed or not open enough for such grasps. Additionally, many detected edges are fake edges that result from partial observations from the camera rather than true graspable edges. For FlowBot3D, we find its performance to be reasonable ($0.57$) when provided with the segmentation mask of the target link (door or drawer) to open, and much lower ($0.2$) without such masks. Note that ArticuBot does not use a segmentation mask for the target link. To form a fair comparison, we also evaluated our policy’s performance after grasping. In such a case, the performance of ArticuBot further improved from $0.75$ to be $0.86$ (See Table III), outperforming FlowBot3D by a large margin.
@@ VI-A Setups
For robust sim2real transfer, we generate more demonstrations in simulation with augmentations on the point cloud observations to make the policy robust to noisy point clouds obtained from real-world depth sensors. Specifically, we add the following two augmentations to the depth map in simulation [9]: the first is edge artifacts that models the noise along object edges, and the second is random holes in the depth map to model random depth pixel value loss in real-world depth cameras. Details of these augmentations can be found in the Appendix -H. We also randomize the camera poses closer to where they are located in the real world. Combined with the non-augmented demonstrations, we generated in total 42.3k trajectories with 4.7M observation-action pairs, and trained a single weighted displacement model high-level policy on this dataset. We find the low-level policy to transfer well without needing to be trained on such point cloud augmentations. We detail the 3 different robot setups as below, visualized in Fig. 3.
@@
We perform camera-to-robot base calibration in all settings, and all point cloud observations are transformed into the robot’s base frame. We use GroundingDino [30] and EfficientSAM [57] to segment the object given a text of the object name, and obtain an object-only point cloud in the real world. For removing the robot from the point cloud, we first render a canonical robot point cloud from the robot urdf and mesh files, transform it to the current point cloud observation using the robot’s current joint angles, and then project it to the 2D depth image using known camera extrinsics and intrinsics to obtain a robot mask. All pixels within the robot mask are removed. We perform additional radius and statistical outlier removing to remove some remaining outlier points from the noisy depth cameras. More details of the real-world perception pipeline can be found in Appendix -I.
@@ VI-B Table-Top Franka Arm Results
The results for all test objects and compared methods in lab A are shown in Fig. 5; the results of ArticuBot in lab B are shown in Table IV. ArticuBot achieves an average grasping success rate of $0.78$ and $0.85$, and a normalized opening performance of $0.63$ and $0.59$ in Lab A and B, respectively, showing that it can generalize to open diverse real articulated objects with varying geometries, shapes and articulation types across both lab settings. We note that some of the test objects are quite challenging and require very precise manipulation, e.g., the knobs of object A2 and A8 are very small, with a diameter of only 2 cm, but ArticuBot can still precisely grasp and open it. As in simulation, we find the grasping success rate of AO-grasp to be low, where it tends to grasp at the “fake” edge in the point cloud due to partial observations, which are not actually graspable (See Appendix -F for visuals of the grasps produced by AO-Grasp). FlowBot3D achieves a reasonable normalized opening performance of $0.38$, starting from the state where the robot gripper already grasps the object. The performance is still lower than ArticuBot, even though ArticuBot performs the additional grasping step. The major failure case for FlowBot3D is that the predicted flow is in the wrong direction, e.g., it predicts upwards flows for opening a microwave (See Appendix -F for visuals of the flows). If we compute the normalized opening performance for ArticuBot only in cases where the grasp is successful (i.e., the same starting conditions as FlowBot3D), the performance of ArticuBot further improves to $0.81$ and outperforms FlowBot3D by a large margin (See Fig. 5). We also find OpenVLA fails to grasp or open any test objects, resulting in a grasping success rate and normalized opening performance of 0. This is likely due to its training data lacking sufficient demonstrations of articulated object manipulation, making generalization to our test cases difficult.
@@
Fig. 6 (zoom-in for better views) visualizes ArticuBot’s predictions on some of the real-world test objects. As shown, the learned per-point weights from the weighted displacement model concentrates on the handle of the object before grasping, and the predicted grasping end-effector pose is quite accurate on the handle, even though for most of the objects the handles are just a very small portion of the point cloud. We note that there is no explicit supervision for the model to learn to assign high weights to the handle; this is automatically learned by just minimizing the imitation learning loss. After grasping, the weights are more randomly distributed across the objects; but as shown in Fig. 6, ArticuBot generalizes to predict different opening end-effector poses for objects with different articulations (left opening revolute joints, right opening revolute joints, pulling out prismatic joints). We do notice a drop in performance compared to the results in simulation. We believe this is mostly due to the noisier point cloud observation from the depth sensors in the real world (e.g., see the noisy point cloud for the green cabinet obtained from the RealSense cameras in Fig. 6). See Fig. LABEL:fig:draw for screenshots of the opening trajectories for more objects (and Appendix -A for trajectories of all objects); please refer to the supplementary materials for videos. Common failure cases for table-top experiments include: 1. The robot arm runs into joint limits while opening the object, due to the limited space of the robot workstation. 2. Wrong end-effector pose predictions for grasping the handle, which we find to happen more for objects with small handles, e.g., A2. Appendix -J provides visualizations of some of the failure cases.
@@ VI-C Mobile X-Arm Results
We do notice a drop in the normalized opening performance compared to the table-top Franka experiments. In our early experiments, we find that Unlike the Franka Arm, the X-Arm lacks impedance control and force sensing. This requires a more precise prediction for the opening end-effector pose. Small prediction errors, such as turning too sharply when opening a revolute door, would result in excessive force for the X-Arm and causes it to stop for motor protection, which is a common failure case in this setting. To partially mitigate this issue, we use the UMI [7] gripper in latter experiments, whose soft, compliant design enhances safety and partially helps prevent the arm from stopping due to excessive force. Another sim2real gap is that, in simulation, objects are isolated by itself; in real kitchens and lounges, objects are usually surrounded by other objects, which occlude its side and top, and only the front side is observable. This additional occlusion might have caused issues for the policy to transfer as well. Finally, the articulation of some objects are inherently ambiguous to judge from a single point cloud observation, e.g., although many dishwashers have a revolute door that open downwards, the dishwasher C2 in Fig. 4 has a prismatic joint that needs to be pulled out horizontally. ArticuBot tends to make more mistakes on such ambiguous objects. See Appendix -J for visualizations of some of the failure cases of ArticuBot.
@@ VII Limitations
Our system currently has the following limitations: 1) Our weighted displacement policy does not handle multi-modal outputs since it is trained with a regression loss (Eq. (3)). This may create issues for it when working with cabinets with multiple doors and opening only a specific one is desired. 2) The current system does not support opening a user-specified door on a multi-door object, as the policy is not trained to be conditioned on any user input. 3) The policy uses point cloud observations. Although this simplifies sim2real transfer, existing depth sensors in the real world do not work well for transparent or reflective objects, and thus our policy would also fail on such objects in the current form. 4) As mentioned above, the X-Arm itself does not support force sensing and impedance control, which requires more precise policy predictions to avoid excessive force. We think adding a force-torque sensor on the X-Arm to enable impedance control could help alleviate this issue; fine-tuning the policy in the real-world via reinforcement learning or a few demonstrations for more precise sub-goal end-effector pose predictions could also help. 5) Objects in real kitchens and lounges are usually occluded by neighboring objects, and we believe that adding this type of occlusion could further improve the sim2real performance of the policy. We leave addressing these limitations as important future work.

### v2

@@ IV-B Policy Learning with a Hierarchical Policy Representation
We then perform cross attention between the scene point cloud and the current end-effector points with Rotary Position Embedding (RoPE) [44], which generates the updated features for current end-effector points as $\{f^{\text{obs-scene}}_{i}\}_{i=1}^{4}$. We generate the features for the goal end-effector points in the same way as for the current end-effector points: $f^{goal}_{i}=[v^{goal}_{i},\text{MLP}^{goal}(ee^{goal}_{i},\delta_{i},\delta^{\prime}_{i})]$. We perform cross attention between the current end-effector points and the goal end-effector points, also with Rotary Position Embedding (RoPE) [44], which produces another set of updated features for the current end-effector points $\{f^{\text{obs-goal}}_{i}\}_{i=1}^{4}$. The final latent embedding used for diffusion is the concatenation of the above two features: $[f^{\text{obs-scene}}_{1},f^{\text{obs-goal}}_{1},...,f^{\text{obs-scene}}_{4},f^{\text{obs-goal}}_{4}]$. This latent embedding is used as the conditioning for an action generation UNet diffusion head, which takes as input this latent conditioning, the robot state (which includes the 3D position, 6D orientation of the end-effector, and finger width), a noisy version of the action, a denoising time step, and predicts the noise. At test time, we use DDIM [42] as the denoising scheduler to generate the actions.
@@ V-C Comparison of Different High-level policies
Please see Appendix -G for more details about these baselines.
@@
Replacing low-level policy with a motion planner $0.24\pm 0.02$
@@ V-D Ablation Studies
  • Replacing low-level policy with a motion planner: This ablation does not use a low-level policy for moving the robot end-effector. We first predict a goal end-effector pose for grasping using the high-level policy and then use a motion planner to reach it. After grasping, we run the high-level policy again to predict a goal end-effector pose for opening the door. We compute the corresponding joint angles using inverse kinematics and use a joint PD controller to reach it.
More details of these ablations can be found in Appendix -I. We compare to these ablations when training with 200 objects without camera randomizations. The results are shown in Table II. We find that using a Point Transformer, not predicting the per-point weights, or predicting a per-point goal end-effector 6D orientation instead of per-point displacements to the goal end-effector points, all lead to worse performance, supporting the effectiveness of our design choices in ArticuBot. Predicting a per-point 6D orientation and averaging them leads to a large drop in performance because it is difficult to correctly compute the average of multiple 6D orientations; in ArticuBot, we avoid this by representing the goal end-effector pose as a collection of points and averaging the displacement to these points. Replacing the low-level policy with motion planning and an IK controller results in very poor performance. We hypothesis it could be due to two reasons: 1. The high-level policy may predict poses with minor collisions, and motion planning often fails due to the inability to find collision-free paths. 2. During door-opening, the joint PD controller takes the shortest joint-space path to the goal pose, ignoring necessary kinematic constraints (e.g., following an arc to open a revolute door), causing the gripper to slip off. Experimentally, the motion planning failure rate is 17%. Among successful grasps, 97% of the failures are due to the gripper falling off the handle during opening (possibly due to ignoring object kinematic constraints). This shows the importance of using a learned low-level policy.
@@ V-E Comparison with Prior Articulated Object Manipulation Methods
The results are shown in Table III, tested without camera randomizations. The grasping success rate of AO-Grasp is much lower than ArticuBot. We find that AO-Grasp often proposes grasps at the edges of the point cloud (e.g., the side wall of a drawer; see Appendix -H for visuals). This likely stems from its training data, which includes many objects in a partially opened state where edge grasps are valid. However, in our test cases, objects are usually closed or not open enough for such grasps. Additionally, many detected edges are fake edges that result from partial observations from the camera rather than true graspable edges. For FlowBot3D, we find its performance to be reasonable ($0.57$) when provided with the segmentation mask of the target link (door or drawer) to open, and much lower ($0.2$) without such masks. Note that ArticuBot does not use a segmentation mask for the target link. To form a fair comparison, we also evaluated our policy’s performance after grasping. In such a case, the performance of ArticuBot further improved from $0.75$ to be $0.86$ (See Table III), outperforming FlowBot3D by a large margin.
@@ V-F Comparison of Different Low-level Policies
V-G Additional Experiments and Evaluations
We show some preliminary experiments in Appendix -E that our hierarchical policy learning approach works on more manipulation tasks beyond articulated object manipulation. We also study how robust the policy is to the orientation and position of the handles, with the results shown in Appendix -F.
@@ VI-A Setups
For robust sim2real transfer, we generate more demonstrations in simulation with augmentations on the point cloud observations to make the policy robust to noisy point clouds obtained from real-world depth sensors. Specifically, we add the following two augmentations to the depth map in simulation [8]: the first is edge artifacts that models the noise along object edges, and the second is random holes in the depth map to model random depth pixel value loss in real-world depth cameras. Details of these augmentations can be found in the Appendix -J. We also randomize the camera poses closer to where they are located in the real world. Combined with the non-augmented demonstrations, we generated in total 42.3k trajectories with 4.7M observation-action pairs, and trained a single weighted displacement model high-level policy on this dataset. We find the low-level policy to transfer well without needing to be trained on such point cloud augmentations. We detail the 3 different robot setups as below, visualized in Fig. 3.
@@
We perform camera-to-robot base calibration in all settings, and all point cloud observations are transformed into the robot’s base frame. We use GroundingDino [29] and EfficientSAM [57] to segment the object given a text of the object name, and obtain an object-only point cloud in the real world. For removing the robot from the point cloud, we first render a canonical robot point cloud from the robot urdf and mesh files, transform it to the current point cloud observation using the robot’s current joint angles, and then project it to the 2D depth image using known camera extrinsics and intrinsics to obtain a robot mask. All pixels within the robot mask are removed. We perform additional radius and statistical outlier removing to remove some remaining outlier points from the noisy depth cameras. More details of the real-world perception pipeline can be found in Appendix -K.
@@ VI-B Table-Top Franka Arm Results
The results for all test objects and compared methods in lab A are shown in Fig. 5; the results of ArticuBot in lab B are shown in Table IV. ArticuBot achieves an average grasping success rate of $0.78$ and $0.85$, and a normalized opening performance of $0.63$ and $0.59$ in Lab A and B, respectively, showing that it can generalize to open diverse real articulated objects with varying geometries, shapes and articulation types across both lab settings. We note that some of the test objects are quite challenging and require very precise manipulation, e.g., the knobs of object A2 and A8 are very small, with a diameter of only 2 cm, but ArticuBot can still precisely grasp and open it. As in simulation, we find the grasping success rate of AO-grasp to be low, where it tends to grasp at the “fake” edge in the point cloud due to partial observations, which are not actually graspable (See Appendix -H for visuals of the grasps produced by AO-Grasp). FlowBot3D achieves a reasonable normalized opening performance of $0.38$, starting from the state where the robot gripper already grasps the object. The performance is still lower than ArticuBot, even though ArticuBot performs the additional grasping step. The major failure case for FlowBot3D is that the predicted flow is in the wrong direction, e.g., it predicts upwards flows for opening a microwave (See Appendix -H for visuals of the flows). If we compute the normalized opening performance for ArticuBot only in cases where the grasp is successful (i.e., the same starting conditions as FlowBot3D), the performance of ArticuBot further improves to $0.81$ and outperforms FlowBot3D by a large margin (See Fig. 5). We also find OpenVLA fails to grasp or open any test objects, resulting in a grasping success rate and normalized opening performance of 0. This is likely due to its training data lacking sufficient demonstrations of articulated object manipulation, making generalization to our test cases difficult.
@@
Fig. 6 (zoom-in for better views) visualizes ArticuBot’s predictions on some of the real-world test objects. As shown, the learned per-point weights from the weighted displacement model concentrates on the handle of the object before grasping, and the predicted grasping end-effector pose is quite accurate on the handle, even though for most of the objects the handles are just a very small portion of the point cloud. We note that there is no explicit supervision for the model to learn to assign high weights to the handle; this is automatically learned by just minimizing the imitation learning loss. After grasping, the weights are more randomly distributed across the objects; but as shown in Fig. 6, ArticuBot generalizes to predict different opening end-effector poses for objects with different articulations (left opening revolute joints, right opening revolute joints, pulling out prismatic joints). We do notice a drop in performance compared to the results in simulation. We believe this is mostly due to the noisier point cloud observation from the depth sensors in the real world (e.g., see the noisy point cloud for the green cabinet obtained from the RealSense cameras in Fig. 6). See Fig. LABEL:fig:draw for screenshots of the opening trajectories for more objects (and Appendix -A for trajectories of all objects); please refer to the supplementary materials for videos. Common failure cases for table-top experiments include: 1. The robot arm runs into joint limits while opening the object, due to the limited space of the robot workstation. 2. Wrong end-effector pose predictions for grasping the handle, which we find to happen more for objects with small handles, e.g., A2. Appendix -L provides visualizations of some of the failure cases.
@@ VI-C Mobile X-Arm Results
We do notice a drop in the normalized opening performance compared to the table-top Franka experiments. In our early experiments, we find that Unlike the Franka Arm, the X-Arm lacks impedance control and force sensing. This requires a more precise prediction for the opening end-effector pose. Small prediction errors, such as turning too sharply when opening a revolute door, would result in excessive force for the X-Arm and causes it to stop for motor protection, which is a common failure case in this setting. To partially mitigate this issue, we use the Fast-UMI [54] gripper in latter experiments, whose soft, compliant design enhances safety and partially helps prevent the arm from stopping due to excessive force. Another sim2real gap is that, in simulation, objects are isolated by itself; in real kitchens and lounges, objects are usually surrounded by other objects, which occlude its side and top, and only the front side is observable. This additional occlusion might have caused issues for the policy to transfer as well. Finally, the articulation of some objects are inherently ambiguous to judge from a single point cloud observation, e.g., although many dishwashers have a revolute door that open downwards, the dishwasher C2 in Fig. 4 has a prismatic joint that needs to be pulled out horizontally. ArticuBot tends to make more mistakes on such ambiguous objects. See Appendix -L for visualizations of some of the failure cases of ArticuBot, and some basic failure recovery abilities of ArticuBot.
@@ VII Limitations
Our system currently has the following limitations: 1) Our weighted displacement policy does not handle multi-modal outputs since it is trained with a regression loss (Eq. (3)). This may create issues for it when working with cabinets with multiple doors and opening only a specific one is desired. 2) The current system does not support opening a user-specified door on a multi-door object, as the policy is not trained to be conditioned on any user input. Although our training data includes multi-door objects, demonstrations are generated for opening the closest door to the initial pose of the robot. The policy learns to open the closest door implicitly, rather than opening a user specified door. 3) The policy uses point cloud observations. Although this simplifies sim2real transfer, existing depth sensors in the real world do not work well for transparent or reflective objects, and thus our policy would also fail on such objects in the current form. 4) As mentioned above, the X-Arm itself does not support force sensing and impedance control, which requires more precise policy predictions to avoid excessive force. We think adding a force-torque sensor on the X-Arm to enable impedance control could help alleviate this issue; fine-tuning the policy in the real-world via reinforcement learning or a few demonstrations for more precise sub-goal end-effector pose predictions could also help. 5) Objects in real kitchens and lounges are usually occluded by neighboring objects, and we believe that adding this type of occlusion could further improve the sim2real performance of the policy. 6) ArticuBot does not use interaction history during the manipulation process. We think that incorporating interaction history with current visual observations could further improve performance, especially for objects whose articulation are ambiguous to judge just from visual observations. We leave addressing these limitations as important future work.
@@ VIII Conclusion
Acknowledgments
This material is based upon work supported by the Toyota Research Institute, National Science Foundation under NSF CAREER Grant No. IIS-2046491, and NIST under Grant No. 70NANB24H314. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of Toyota Research Institute, National Science Foundation, or NIST.